OpenStudy (anonymous):

y over negative 2+ 5 = 12

OpenStudy (anonymous):

y=36

OpenStudy (anonymous):

no way the answer choices are −14 14 34 −34

OpenStudy (anonymous):

Is ur question y/(-2+5)=12?

OpenStudy (anonymous):

y over negative 2+ 5 = 12

OpenStudy (anonymous):

-y/2+5=12 (-y+10)=24 -y=14 y=-14